There is a circular track of 440 meters. There are three persons Symonds, Harbhajan and Inzmam standing on different points on the track ready to start the race. Symonds and Harbhajan are standing diametrically opposite to each other while Inzmam is midway between them such that Symonds, Inzmam and Harbhajan are standing in clockwise direction. The race started at 10 AM. The speed of Symonds, Harbhajan and Inzmam are respectively 5 m/s, 10 m/s and 8 m/s. At what time would Harbhajan and Inzmam meet for the second time if all of three run in clockwise direction ?
Answer by ankor@dixie-net.com(22740)

There is a circular track of 440 meters.
There are three persons Symonds, Harbhajan and Inzmam standing on different points on the track ready to start the race.
Symonds and Harbhajan are standing diametrically opposite to each other while Inzmam is midway between them such that Symonds, Inzmam and Harbhajan are standing in clockwise direction.
The race started at 10 AM. The speed of Symonds, Harbhajan and Inzmam are respectively 5 m/s, 10 m/s and 8 m/s.
At what time would Harbhajan and Inzmam meet for the second time if all of three run in clockwise direction ?
:
Looks like Symon is irrelevant to this problem
:
The relative speed between Harb and Inz: 10 - 8 = 2 m/s and they are 330m apart
Therefore, 330/2 = 165 sec; 2 min 45 sec; when they meet the 1st time it will be 10:02:45
When they meet the 2nd time: 440/2 = 220 sec; 3 min 40 sec
10:02:45
+ 03:40
--------------
10:06:25 when they meet the 2nd time